# neurips_2025.bib Edge-ASR: Towards Low-Bit Quantization of Automatic Speech Recognition Models

### v1

@@ Abstract
Recent advances in Automatic Speech Recognition (ASR) have demonstrated remarkable accuracy and robustness in diverse audio applications, such as live transcription and voice command processing. However, deploying these models on resource-constrained edge devices (e.g., IoT device, wearables) still presents substantial challenges due to strict limits on memory, compute and power. Quantization, particularly Post-Training Quantization (PTQ), offers an effective way to reduce model size and inference cost without retraining. Despite its importance, the performance implications of various advanced quantization methods and bit-width configurations on ASR models remain unclear. In this work, we present a comprehensive benchmark of eight state-of-the-art (SOTA) PTQ methods applied to two leading edge-ASR model families, Whisper and Moonshine. We systematically evaluate model performances (i.e., accuracy, memory I/O and bit operations) across seven diverse datasets from the open ASR leaderboard, analyzing the impact of quantization and various configurations on both weights and activations. Built on an extension of the LLM compression toolkit, our framework integrates edge-ASR models, diverse advanced quantization algorithms, a unified calibration and evaluation data pipeline, and detailed analysis tools. Our results characterize the trade-offs between efficiency and accuracy, demonstrating that even -bit quantization can succeed on high capacity models when using advanced PTQ techniques. These findings provide valuable insights for optimizing ASR models on low-power, always-on edge devices.
@@ 1 Introduction
Real-time Automatic Speech Recognition (ASR) models have revolutionized the natural audio processing pipeline and achieved remarkable accuracy and robustness in diverse applications such as live transcription, voice command recognition and assistive use cases ([ASRSurvey]). As interest in on-device ASR continues to grow, there is an emerging demand for models that operate efficiently on low-power, always-on edge hardware (e.g., IoT device, wearables). Among the state-of-the-art (SOTA) edge-ASR models, OpenAI’s Whisper ([Whisper]) has demonstrated strong accuracy in compact model sizes, making it a practical candidate for edge inference. Meanwhile, Useful Sensors’ Moonshine ([Moonshine]) reduces latency and adapts well to variable-length audio inputs, further optimized for resource-constrained deployment. Both model families currently lead the open ASR leaderboard ([Leaderboard]) across seven evaluation datasets, among model sizes ranging from M to M parameters.
@@
While many advanced PTQ techniques have been proposed and studied in the context of large language models (LLMs) ([QuantLLM]), their applications to edge-ASR remain unexplored. Exisiting quantization toolkits and benchmarks ([LLMC]) focus on LLM model architectures, and task-specific datasets that do not apply to speech recognition directly. Furthermore, although the open ASR leaderboard provides standardized full precision baselines, it lacks coverage of quantized models, leaving the performance trade-offs from quantization in this domain largely uncharacterized.
In this work, we fill this gap by benchmarking eight PTQ methods applied to two Transformer-based edge-ASR families, Whisper and Moonshine. The quantization methods are categorized into three primary strategies: transformation-based, reconstruction-based and rounding optimization. We systematically evaluate quantization of weights and activations across seven public ASR datasets, measuring accuracy, memory I/O and bit operations (BOPs) ([BOP]). We also perform ablation studies on calibration data sources and sizes, and analyze quantization sensitivity. All experiments are built on an extension of the LLM compression toolkit ([LLMC]), enabling a reproducible and extensible ASR quantization framework. The overall workflow is demonstrated in Figure 1.
Contributions. (a) We deliver the first systematic benchmark of quantized edge-ASR models, assessing eight advanced PTQ methods over multiple datasets, architectures and configurations, to establish a quantized standard for the open ASR leaderboard. (b) We extend and contribute to the LLM compression toolkit to support ASR models, enabling an end-to-end, extensible quantization workflow for speech recognition. (c) Through extensive evaluations, analysis and comprehensive comparisons from various factors (e.g., model accuracy, memory I/O, BOPs), we provide valuable insights into how different quantization methods and configurations impact model performance, offering guidance for deploying ASR on memory and compute constrained low-power edge devices.
@@ 2.1 Quantization Algorithms
Post-Training Quantization algorithms can be broadly categorized into three primary strategies: scaling-based transformation, reconstruction-based optimization, and rounding-based optimization. Scaling-based transformation techniques, such as SmoothQuant ([SmoothQuant]), AWQ ([AWQ]), and OmniQuant ([OmniQuant]), apply learned or heuristic scaling to weights and activations, reshaping their distributions to be more quantization-friendly. Reconstruction-based methods, such as GPTQ ([GPTQ]), reconstruct block outputs by minimizing quantization error iteratively, preserving the original model behavior after quantization. Rounding-based optimizations refine the rounding process to better preserve accuracy. Representitive methods are naïve round-to-nearest (RTN) and TesseraQ ([TesseraQ]), where an iterative and progressive adaptive rounding is proposed. In addition, methods like QUIK ([QUIK]) and SpQR ([SpQR]) use hybrid quantization strategies to balanace model accuracy and overall compression ratio. Table 1 summarizes the eight evaluated quantization algorithms.
@@
To evaluate the impact of post-training quantization techniques on edge-ASR performance, we select two state-of-the-art open-source ASR model families: OpenAI’s Whisper ([Whisper]) and Useful Sensors’ Moonshine ([Moonshine]). Both model families lead the open ASR leaderboard across seven diverse datasets and feature compact model sizes (M to M parameters), making them well-suited for deployment on low-power, always-on edge devices.
@@ 2.3 Benchmarks and Datasets
Following the open ASR leaderboard, we benchmark eight PTQ methods on two edge-ASR model families across seven evaluation datasets: AMI corpus ([AMI]), Eearnings-22 ([Earnings-22]), GigaSpeech ([GigaSpeech]), LibriSpeech ([Librispeech]), SPGISpeech ([SPGISpeech]), TED-Lium ([TED-LIUM]) and VoxPopuli ([VoxPopuli]). These datasets cover a diverse range of domains, accents and acoustic conditions, providing a robust performance baseline.
For each quantization method, we explore multiple configurations: varying weight and activation bit-widths and quantization granularities. To align with typical support in low-power inference accelerators, we evaluate per-channel and per-group quantization for weights, and per-tensor and per-token quantization for activations. Our evaluation reports a full spectrum of deployment-relevant metrics, including word-error-rate (WER), model size, memory I/O and BOPs. The comparisons offer end-to-end insights into the trade-offs between accuracy and efficiency under different quantization strategies.
@@ 3 Benchmarking Quantized edge-ASR Models
This section presents a comprehensive benchmarking of PTQ methods on edge-ASR models. We begin with the experimental setup (subsection 3.1), followed by an in-depth quantization result and analysis (subsection 3.2), covering bit-width selection, ultra-low bit quantization, granularity, weight clipping, and the use of symmetric versus asymmetric quantization. We then compare deployment-relevant metrics, including WER, model size, memory I/O and BOPs in subsection 3.3. Additional ablation studies on the impact of calibration data source and size are further discussed in Appendix E.
@@ 3.1 Experimental Setup
Models. We assess two ASR families optimized for edge deployment. Three variants: Tiny (M), Base (M) and Small (M), are selected from the Whisper family, while the Moonshine family includes Tiny (M) and Base (M). Both model families share an encoder-decoder Transformer architecture but differ in positional embedding and configurations. We employ a byte-level BPE tokenizer (vocabularies of for Whisper and for Moonshine), and apply standard text normalization before computing WER.
Datasets. Following the open ASR leaderboad, we evaluate on seven benchmark datasets. These datasets cover diverse domains, accents and speaking styles, with audio durations ranging from to hours. Our floating-point baselines match reported leaderboard results ([Leaderboard]). For algorithms that require calibration, we use a fixed set of English utterances sampled from the Mozilla Common Voice ([CommonVoice]) that are different from any evaluation dataset, ensuring a fair comparison. Table 2 summarizes the full benchmark setup.
@@
| Datasets | AMI corpus, Eearnings-22, GigaSpeech, LibriSpeech (clean, other), SPGISpeech, TED-Lium, VoxPopuli |
@@ 3.2.1 Bit-width selection
To determine the optimal bit-width for ASR models, we evaluate WER under various PTQ algroithms across all benchmark datasets. Table 3 reports the averaged WER of the quantized edge-ASR models across datasets, alongside its floating-point baseline. Detailed per-dataset WER performances of each model are provided in Appendix D.
Our results show that PTQ with -bit weights and -bit activations (w8-a8) reliably preserves model performance across all models and quantization algorithms. Within both model families, reducing weight precision to bits leads to notable degradation in smaller models such as Whisper Tiny (Table 11) and Moonshine Tiny (Table 14), while larger models, including Whisper Base (Table 12), Whisper Small (Table 13) and Moonshine Base (Table 15) exhibit stronger robustness under the same low-bit quantization. For the Moonshine family, SmoothQuant performs poorly at low weight bit-width, and only becomes effective when weight precision increases to -bit.
@@ 3.2.2 Ultra-low bit quantization
Enforcing -bit weight precision causes all PTQ methods to fail, resulting unacceptably high WER. For weights below -bit, Quantization-Awared Training (QAT) becomes essential. However, at -bit weight precision, several algorithms produce promising results. As shown in Table 4, OmniQuant and hybrid strategies (e.g., QUIK, SpQR that preserves higher precision for the top outliers) maintain reasonable accuracy. In contrast, scaling-based transformation such as SmoothQuant perform poorly on the Whisper Base model. Similarly, RTN and TesseraQ struggle on the Moonshine Base model at this precision, indicating that their rounding-based strategies alone are insufficient at ultra-low precision.
To further examine each algorithm’s internal behavior, we analyze the Whisper Base model and present the averaged kurtosis value of weights and activations over all layers in Figure 2. This provides a statistical perspective on the distribution characteristics after quantization. We observed that scaling-based transformation methods (SmoothQuant, AWQ) achieve lower kurtosis for activations but at the cost of increased weight kurtosis, due to the design principle of migrating the quantization difficulty from activations to weights. This explains significant performance degradation under ultra-low bit settings, even when activations remain at bits. These results underscore the need for algorithmic roubustness and outlier handling when pushing PTQ to its limits.
@@ 3.2.3 Quantization granularity
To assess the impact of weight and activation quantization granulairty, we select three representitive PTQ algorithms (AWQ, GPTQ and TesseraQ), one from each categrory, and apply -bit uniform quantization to weights under two schemes: per-channel and per-group. -bit per-tensor and per-token quantization are applied on activations. Table 5 reports the averaged WER across all seven datasets for each model and granularity setting.
@@
Our results show that GPTQ consistently delivers the lowest averaged WER across all weight granularity configurations, demonstrating the strong robustness to coarse quantization. For larger models (Moonshine Base and Whisper Small), per-channel weight quantization shows reasonable accuracy, making it an efficient option. In contrast, smaller models (Whisper Tiny and Whisper Base) benefit notably from per-group quantization, where finer-grained scaling helps to mitigate quantization noise and significantly improves WER. We also observe that activation granulairty has minimal impact at -bit precision. These findings highlight the importance of choosing appropriate weight quantization granularity, especially when targeting low capacity models on hardwares with limited quantization support.
@@ 3.2.4 Weight clipping
The technique of weight clipping is commonly used to constrain the range of weight values prior to quantization. By suppressing extreme outliers, it reduces quantization error. However, this process may also introduce clipping loss, potentially harming model accuracy. To investigate this trade-off, we use AWQ as a case study to evaluate the impact of weight clipping under four bit-width settings: w4-a8, w4-a16, w8-a8 and w8-a16 on both Whisper Tiny and Whisper Base models. Table 6 reports the resulting averaged WER across all seven evaluation datasets.
@@
Our experiments reveal a clear capacity-dependent effect. For Whisper Tiny model, weight clipping consistently degrades performance across all bit-width settings. The clipping error introduced by capping weight values outweighs any reduction in quantization noise, leading to higher WER. For Whisper Base model, weight clipping begins to show measurable benefits as precision increases. At -bit weight settings, clipped models achieve modest WER reductions compared to unclipped counterparts. These findings indicate that weight clipping should be applied selectively. It can enhance quantization robustness for higher capacity models operating at higher precisions, but tend to harm smaller models, which are less resilient to additional clipping distortion.
@@ 3.2.5 Symmetric or asymmetric
Symmetric quantization with a zero offset is generally more hardware-friendly and thus preferred for deployment. However, asymmetric quantization introduces an additional grid point in the quantized space, which can improve model accuracy, especially under low bit-width settings. To quantify this effect, we compare symmetric and asymmetric weight quantization across all eight PTQ algorithms on Whisper Base and Moonshine Base models. As shown in Table 7, asymmetric quantization outperforms symmetric quantization in out of algorithms. Notably, for algorithms like SmoothQuant and AWQ, asymmetric quantization proves to be crucial to achieving acceptable WER when using -bit per-channel weight quantization. These results hightlight that, despite its slightly higher implementation complexity, asymmetric quantization often yields superior performance, making it a valuable strategy in low-bit deployment scenarios.

### v2

# Edge-ASR: Towards Low-Bit Quantization of Automatic Speech Recognition Models

@@ Abstract
Recent advances in Automatic Speech Recognition (ASR) have demonstrated remarkable accuracy and robustness in diverse audio applications, such as live transcription and voice command processing. However, deploying these models on resource-constrained edge devices (e.g., IoT device, wearables) still presents substantial challenges due to strict limits on memory, compute and power. Quantization, particularly Post-Training Quantization (PTQ), offers an effective way to reduce model size and inference cost without retraining. Despite its importance, the performance implications of various advanced quantization methods and bit-width configurations on ASR models remain unclear. In this work, we present a comprehensive benchmark of eight state-of-the-art (SOTA) PTQ methods applied to two leading edge-ASR model families, Whisper and Moonshine. We systematically evaluate model performances (i.e., accuracy, memory I/O and bit operations) across seven diverse datasets from the open ASR leader-board, analyzing the impact of quantization and various configurations on both weights and activations. Built on an extension of the LLM compression toolkit, our framework integrates edge-ASR models, diverse advanced quantization algorithms, a unified calibration and evaluation data pipeline, with detailed analysis tools. Our results characterize the trade-offs between efficiency and accuracy, demonstrating that even -bit quantization can succeed on high capacity models when using advanced PTQ techniques. These findings provide valuable insights for optimizing ASR models on low-power, always-on edge devices.
@@ 1 Introduction
Real-time Automatic Speech Recognition (ASR) models have revolutionized the natural audio processing pipeline and achieved remarkable accuracy and robustness in diverse applications such as live transcription, voice command recognition and assistive use cases ([11]). As interest in on-device ASR continues to grow, there is an emerging demand for models that operate efficiently on low-power, always-on edge hardware (e.g., IoT device, wearable). Among the state-of-the-art (SOTA) edge-ASR models, OpenAI’s Whisper ([17]) has demonstrated strong accuracy in compact model sizes, making it a practical candidate for edge inference. Meanwhile, Useful Sensors’ Moonshine ([10]) reduces latency and adapts well to variable-length audio inputs, which is further optimized for resource-constrained deployment. Both model families currently lead the open ASR leader-board ([3]) across seven evaluation datasets, among model sizes ranging from M to M parameters.
@@
While many advanced PTQ techniques have been proposed and studied in the context of large language models (LLMs) ([12]), their applications to edge-ASR remain unexplored. Existing quantization toolkits and benchmarks ([8]) focus on LLM model architectures, and task-specific datasets that do not apply to speech recognition directly. Furthermore, although the open ASR leader-board provides standardized full precision baselines, it lacks coverage of quantized models, leaving the performance trade-offs from quantization in this domain largely unexplored.
In this work, we fill this gap by benchmarking eight PTQ methods applied to two Transformer-based edge-ASR families, Whisper and Moonshine. The quantization methods are categorized into three primary strategies: transformation-based, reconstruction-based and rounding optimization. We systematically evaluate quantization of weights and activations across seven public ASR datasets, measuring accuracy, memory I/O and bit operations (BOPs) ([4]). We also perform ablation studies on calibration data sources and sizes, and analyze quantization sensitivity. All experiments are built on an extension of the LLM compression toolkit ([8]), enabling a reproducible and extensible ASR quantization framework. The overall workflow is demonstrated in Figure 1.
Contributions. (a) We deliver the first systematic benchmark of quantized edge-ASR models, assessing eight advanced PTQ methods over multiple datasets, architectures and configurations, to establish a quantized standard for the open ASR leader-board. (b) We extend and contribute to the LLM compression toolkit to support ASR models, enabling an end-to-end, extensible quantization workflow for speech recognition. (c) Through extensive evaluations, analysis and comprehensive comparisons from various factors (e.g., model accuracy, memory I/O, BOPs), we provide valuable insights into how different quantization methods and configurations impact model performance, offering guidance for deploying ASR on memory and compute constrained low-power edge devices.
@@ 2.1 Quantization Algorithms
Post-Training Quantization algorithms can be broadly categorized into three primary strategies: scaling-based transformation, reconstruction-based optimization, and rounding-based optimization. Scaling-based transformation techniques, such as SmoothQuant ([22]), AWQ ([14]), and OmniQuant ([20]), apply learned or heuristic scaling to weights and activations, reshaping their distributions to be more quantization-friendly. Reconstruction-based methods, such as GPTQ ([7]), reconstruct block outputs by minimizing quantization error iteratively, preserving the original model behavior after quantization. Rounding-based optimizations refine the rounding process to better preserve accuracy. Representative methods are naïve round-to-nearest (RTN) and TesseraQ ([13]), where an iterative and progressive adaptive rounding is proposed. In addition, methods like QUIK ([2]) and SpQR ([6]) use hybrid quantization strategies to balance model accuracy and overall compression ratio. Table 1 summarizes the eight evaluated quantization algorithms.
@@ 2.2 Edge-ASR Models
To evaluate the impact of post-training quantization techniques on edge-ASR performance, we select two state-of-the-art open-source ASR model families: OpenAI’s Whisper ([17]) and Useful Sensors’ Moonshine ([10]). Both model families lead the open ASR leader-board across seven diverse datasets and feature compact model sizes (M to M parameters), making them well-suited for deployment on low-power, always-on edge devices.
@@ 2.3 Benchmarks and Datasets
Following the open ASR leaderboard, we benchmark eight PTQ methods on two edge-ASR model families across seven evaluation datasets: AMI corpus ([19]), Earnings-22 ([18]), GigaSpeech ([5]), LibriSpeech ([16]), SPGISpeech ([15]), TED-Lium ([9]) and VoxPopuli ([21]). These datasets cover a diverse range of domains, accents and acoustic conditions, providing a robust performance baseline.
For each quantization method, we explore multiple configurations: varying weight and activation bit-widths and quantization granularity. To align with typical support in low-power inference accelerators, we evaluate per-channel and per-group quantization for weights, and per-tensor and per-token quantization for activations. Our evaluation reports a full spectrum of deployment-relevant metrics, including word-error-rate (WER), model size, memory I/O and BOPs. The comparisons offer end-to-end insights into the trade-offs between accuracy and efficiency under different quantization strategies.
@@ 3 Benchmarking Quantized edge-ASR Models
This section presents a comprehensive benchmarking of PTQ methods on edge-ASR models. We begin with the experimental setup (subsection 3.1), followed by an in-depth quantization result analysis (subsection 3.2), covering bit-width selection, ultra-low bit quantization, granularity, weight clipping, and the use of symmetric versus asymmetric quantization. We also compare deployment relevant metrics, including WER, model size, memory I/O and BOPs in subsection 3.3. Additional ablation studies on the impact of calibration data source and size are further discussed in Appendix E.
@@ 3.1 Experimental Setup
Models. We assess two ASR families optimized for edge deployment. Three variants: Tiny (M), Base (M) and Small (M), are selected from the Whisper family, while the Moonshine family includes Tiny (M) and Base (M). Both model families share an encoder-decoder Transformer architecture but differ in positional embedding and other configurations. We employ a byte-level BPE tokenizer (vocabularies of for Whisper and for Moonshine), and apply standard text normalization before computing WER.
Datasets. Following the open ASR leader-board, we evaluate on seven benchmark datasets. These datasets cover diverse domains, accents and speaking styles, with audio durations ranging from to hours. Our floating-point baselines match reported leader-board results ([3]). For algorithms that require calibration, we use a fixed set of English utterances sampled from the Mozilla Common Voice ([1]) that are different from any evaluation dataset, ensuring a fair comparison. Table 2 summarizes the full benchmark setup.
@@
| Datasets | AMI corpus, Earnings-22, GigaSpeech, LibriSpeech (clean, other), SPGISpeech, TED-Lium, VoxPopuli |
@@ 3.2.1 Bit-width selection
To determine the optimal bit-width for ASR models, we evaluate WER under various PTQ algorithms across all benchmark datasets. Table 3 reports the averaged WER of the quantized edge-ASR models across datasets, alongside its floating-point baseline. Detailed per-dataset WER performances of each model are provided in Appendix D.
Our results show that PTQ with -bit weights and -bit activations (w8-a8) reliably preserves model performance across all models and quantization algorithms. Within both model families, reducing weight precision to bits leads to notable degradation of performance metrics in smaller models such as Whisper Tiny (Table 11) and Moonshine Tiny (Table 14), while larger models, including Whisper Base (Table 12), Whisper Small (Table 13) and Moonshine Base (Table 15) exhibit strong robustness under the same low-bit quantization. For the Moonshine family, SmoothQuant performs poorly at low weight bit-width, and only becomes effective when weight precision increases to -bit.
@@ 3.2.2 Ultra-low bit quantization
Enforcing -bit weight precision causes all PTQ methods to fail, resulting in high WER. For weights below -bit, Quantization-Aware Training (QAT) becomes essential. However, at -bit weight precision, several algorithms still produce promising results. As shown in Table 4, OmniQuant and hybrid strategies (e.g., QUIK, SpQR that preserves higher precision for the top outliers) maintain reasonable accuracy. In contrast, scaling-based transformation such as SmoothQuant perform poorly on the Whisper Base model. Similarly, RTN and TesseraQ struggle on the Moonshine Base model at this precision, indicating that their rounding-based strategies alone are insufficient at ultra-low precision.
To further examine each algorithm’s internal behavior, we analyze the Whisper Base model and present the averaged kurtosis value of weights and activations over all layers in Figure 2. This provides a statistical perspective on the distribution characteristics after quantization. We observed that scaling-based transformation methods (SmoothQuant, AWQ) achieve lower kurtosis for activations at the cost of increased weight kurtosis, due to the design principle of offloading the quantization difficulty from activations to weights. This justifies the significant performance degradation under ultra-low bit settings, even when activations remain at bits. These results underscore the need for algorithmic robustness and outlier handling when pushing PTQ to its limits.
@@ 3.2.3 Quantization granularity
To assess the impact of weight and activation quantization granularity, we select three representative PTQ algorithms (AWQ, GPTQ and TesseraQ), one from each category, and apply -bit uniform quantization to weights under two schemes: per-channel and per-group. -bit per-tensor and per-token quantization are applied on activations. Table 5 reports the averaged WER across all seven datasets for each model and granularity setting.
@@
Our results show that GPTQ consistently delivers the lowest averaged WER across all weight granularity configurations, demonstrating the strong robustness to coarse quantization. For larger models (Moonshine Base and Whisper Small), per-channel weight quantization shows reasonable accuracy, making it an efficient option. In contrast, smaller models (Whisper Tiny and Whisper Base) benefit notably from per-group quantization, where fine-grained scaling helps mitigate quantization noise and significantly improves WER. We also observe that activation granularity has minimal impact at -bit precision. These findings highlight the importance of choosing the appropriate weight quantization granularity, especially when targeting low capacity models on hardware with limited quantization support.
@@ 3.2.4 Weight clipping
The technique of weight clipping is commonly used to constrain the range of weight values prior to quantization. By suppressing extreme outliers, it reduces quantization error. However, this process may also introduce clipping loss, potentially harming model accuracy. To investigate this trade-off, we use AWQ as a case study to evaluate the impact of weight clipping under four bit-width settings: w4-a8, w4-a16, w8-a8 and w8-a16 on both Whisper Tiny and Whisper Base models. Table 6 reports the resulting average WER across all seven evaluation datasets.
@@
Our experiments reveal a clear capacity-dependent effect. For Whisper Tiny model, weight clipping consistently degrades performance across all bit-width settings. The clipping error introduced by capping weight values outweighs any reduction in quantization noise, leading to higher WER. For Whisper Base model, weight clipping demonstrates measurable benefits as precision increases. At -bit weight settings, clipped models achieve modest WER reductions compared to their unclipped counterparts. These findings indicate that weight clipping should be applied selectively. It can enhance quantization robustness for higher capacity models operating at higher precisions, but tend to harm smaller models, which are less resilient to additional clipping distortion.
@@ 3.2.5 Symmetric or asymmetric
Symmetric quantization with a zero offset is generally more hardware-friendly and thus preferred for deployment. However, asymmetric quantization introduces an additional grid point in the quantized space, which can improve model accuracy, especially under low bit-width settings. To quantify this effect, we compare symmetric and asymmetric weight quantization across all eight PTQ algorithms on Whisper Base and Moonshine Base models. As shown in Table 7, asymmetric quantization outperforms symmetric quantization in out of algorithms. Notably, for algorithms like SmoothQuant and AWQ, asymmetric quantization proves to be crucial to achieving acceptable WER when using -bit per-channel weight quantization. These results highlight that, despite its slightly higher implementation complexity, asymmetric quantization often yields superior performance, making it a valuable strategy in low-bit deployment scenarios.